OpenStudy (angelina__nievse):

Can someone check my work will medal. 1 what is equivalent to (8^2)-4 I choose 1/8.8.8.8.8.8.8.8. Question 2 simplfy 11^-4 /11^8 I choose 1/11^4

OpenStudy (mathmate):

The first one is correct, (8^2)^(-4)=8^(2*(-4)=8^(-8). But 11^(-4)/11^(8) you need the rule of exponents: x^a/x^b=x^(a-b) So 11^(-4)/11^(8)=11^(-4-8)=11^(-12)